OpenStudy (anonymous):

It's given that 2x + y=5 X= -4 How do i get y? I'm so stuck on this please help me

OpenStudy (anonymous):

substitute for x. it tells you the value for it ..

OpenStudy (anonymous):

-8 right?

OpenStudy (anonymous):

then what's next?

OpenStudy (anonymous):

-8 + y = 5

OpenStudy (anonymous):

as per the substitution...

OpenStudy (anonymous):

5-8=y? Y=-3

OpenStudy (anonymous):

is it the next step?

OpenStudy (anonymous):

nope, you add 8 to both sides. what you did there was add 5 to both sides, and 5+5 != 0

OpenStudy (anonymous):

5+8=y y=13

OpenStudy (anonymous):

is it the answer?

OpenStudy (anonymous):



OpenStudy (anonymous):

so yes, @michu you are correct :D

OpenStudy (dwelsher):

1.substitute -4 for x in the original equation 2.-8+y=-4 3.moving the - 8 to the other side of the equation changes the sign to +8 4.+8+(-4)=?=4 5.y=4